In a precipitate reaction between BaCl2(aq) and Na3PO4(aq), 57 mL of 0.100 M BaCl2(aq) completely reacted with 25.7 mL of Na3PO4(aq). What was the molarity of Na3PO4(aq)?

3 BaCl2(aq) + 2 Na3PO4(aq) → Ba3(PO4)2(s) + 6 NaCl(aq)
Answer #1

3 BaCl2(aq) + 2 Na3PO4(aq) → Ba3(PO4)2(s) + 6 NaCl(aq)

using the dilution law M1V1 /n1 = M2V2 /n2

where M1 = 0.100M , V1 = 0.057 L , n1 = 3   ===============> for BaCl2

           M2 = ?         , V2 = 0.0257 L , n2 = 2 =================> for Na3PO4

SUBSTITUTING THE VALUES

M2 = M1*V1*n2 / n1*V2

=> M2 = 0.100*0.057*2 / 3*0.0257

=> M2 = 0.1478M

answer => 0.1478 M of Na3PO4